OpenStudy (anonymous):

find the slope of m=9-5/-7-(-7)

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

undefined

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

calculate slope between (10, 93) and (4, 39)

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

-9

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

(7, -8) and (7, 3)

OpenStudy (anonymous):

I got 0

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

undefined?

jimthompson5910 (jim_thompson5910):

Hint: whenever the x coordinates are the same, you have a vertical line

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

vertical lines have undefined slopes

OpenStudy (anonymous):

(5-2) and (5, -5) =3?

jimthompson5910 (jim_thompson5910):

show me your work on that one

OpenStudy (anonymous):

-2--5=3 or is it 0?

jimthompson5910 (jim_thompson5910):

that part is true

jimthompson5910 (jim_thompson5910):

but you didn't finish up

OpenStudy (anonymous):

5-5=0

jimthompson5910 (jim_thompson5910):

so that means the slope is what

OpenStudy (anonymous):

0?

jimthompson5910 (jim_thompson5910):

slope is change in ____ over change in _____

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

but then the answer must be undefined

jimthompson5910 (jim_thompson5910):

why

OpenStudy (anonymous):

because if -2--5=3, and 0-3=0, it can not be defined

jimthompson5910 (jim_thompson5910):

you mean 5-5 = 0 right?

OpenStudy (anonymous):

sorry, yes

jimthompson5910 (jim_thompson5910):

slope is change in y over change in x

jimthompson5910 (jim_thompson5910):

slope = rise/run = 3/0 = undefined

OpenStudy (anonymous):

(-2,-21) and (4,27)=1/8?

jimthompson5910 (jim_thompson5910):

last one and no

OpenStudy (anonymous):

-8?

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

-1/8?

jimthompson5910 (jim_thompson5910):

you're guessing

OpenStudy (anonymous):

m=(-2, -21) (4, 27) -2-4 and 27-21=8

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (27-(-21))/(4-(-2)) m = (27+21)/(4+2) m = 48/6 m = 8

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I was confusing myself at first

jimthompson5910 (jim_thompson5910):

yw